Exercise 24-9 Partially correct answer. Your answer is partially correct. Try again. Legend Service Center just purchased an automobile hoist for $32,900. The hoist has an 8-year life and an estimated salvage value of $3,640. Installation costs and freight charges were $4,140 and $830, respectively. Legend uses straight-line depreciation. The new hoist will be used to replace mufflers and tires on automobiles. Legend estimates that the new hoist will enable his mechanics to replace 6 extra mufflers per week. Each muffler sells for $73 installed. The cost of a muffler is $38, and the labor cost to install a muffler is $12. (a) Compute the cash payback period for the new hoist. (Round answer to 2 decimal places, e.g. 10.50.) Cash payback period Entry field with correct answer 5.28 years (b) Compute the annual rate of return for the new hoist. (Round answer to 1 decimal place, e.g. 10.5.) Annual rate of return Entry field with incorrect answer 7.65 %
Answer #1

ANSWER:-

CALCULATE NET INCOME & NET CASH FLOW :

SALES (6*73*52) 22776
LESS: COST OF MUFFLER (38*6*52) (11856)
LESS: LABOUR COST (12*6*52) (3744)
NET CASH FLOW 7176
LESS: DEPRECIATION (32900+4140+830-3640)/8 (4278.75)
NET INCOME 2897.75

CASH PAYBACK PERIOD = INITIAL INVESTMENT /ANNUAL CASH FLOW

= (32900+4140+830) / 7176

=37870 / 7176

=5.277

CASH PAYBACK PERIOD = 5.27 YEARS

ACCOUNTING RATE OF RETURN = NET INCOME*100/INITIAL INVESTMENT

= 2897.75*100 / (32900+4140+830-3640)

=2897.75*100 / 34230       

   =2897.75 / 34230

= 8.4%

ACCOUNTING RATE OF RETURN = 8.4%

--------------------------------------

if you have any query please ask me in comment box i am here to helps you don't give direct Thumbs down.if you satisfied my work give Thumbs UP

                                                *******THANK YOU*******